A rectangular piece of land has been claimed using a 160 yard cord. If the length of this rectangular lot is 20 yards more than twice it's width, then what are the dimensions of the lot
Answer by ikleyn(52780)

.
A rectangular piece of land has been claimed using a 160 yard cord. If the length of this rectangular lot
is 20 yards more than twice it's width, then what are the dimensions of the lot
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

If "w" represents the width, then the length is 2w+20.


The perimeter is (2w+20) + w + (2w+20) + w = 6w + 40.


Hence, the equation is 6w + 40 = 160,  --->  6w = 120  --->  w = {{120/6}}} = 20.


Answer.  The width is 20 ft. The length is 2w + 20 = 2*20+20 = 60 ft.

For many other similar solved problems see the lessons
    - Solved problems on the perimeter and side lengths of a triangle
    - Solved problems on the perimeter and side lengths of a right-angled triangle
    - Solved problems on the perimeter and the side lengths of a parallelogram
    - Solved problems on the perimeter and side lengths of a rectangle
in this site.